OpenStudy (anonymous):

The graph of y=3^(x) is shifted right 2 and up 1. Which of the following shows these transformations?

OpenStudy (anonymous):

y-1=3^(x-2)

OpenStudy (anonymous):

A. y=3^(x+2)+1 B. y=3^(x-2)+1 C. y=3^2x+1 D. y=3^(x+2)-1